Adv. Math. Econ. 10, 91-100 (2007)

A remark on law invariant convex risk measures

Shigeo Kusuoka

Graduate School of Mathematical Sciences, The University of Tokyo, Komaba 3-8-1, Meguro-ku, Tokyo 153-8914, Japan Received: September 11, 2006 Revised: November 22, 2006 JEL classification: C65, G19 Mathematical Subject Classification (2000): 60B05 Abstract. The author gives a simple proof of the representation theorem for law invariant convex risk measures which was obtained by Kusuoka(Adv.Math.Econ. 3:83–95,2001), Frittelli and Rossaza Gianin(Adv.Math.Econ. 7:33–46,2005) and Jouini et al.(Adv.Math.Econ. 9:49–71,2006). Key words: risk measure, law invariant